The price of a Boston Red Sox baseball game rises from $10 to $12 a game. The quantity of tickets sold falls from 160,000 tickets to 144,000. b.
Answer #1

Answer :

Price elasticity of demand = % change on demand /% change in price

% change in demand =(144000-160000)/160000=-10%

% change in price =(12-10)/10=20%

Price elasticity of demand =-10%/20%=-0.5

Since price elasticity is less than 1 so it is relatively inelastic demand.